OpenStudy (anonymous):

in 2000 the population of a state was 4 million and growing continously at a rate of 3.3% per year (a). write a function that modela the state's population inmillions x years after 200 (b). graph f in [0,10] by[4,7,1] (c). estimate thje state's populaton in 2008

OpenStudy (anonymous):

a) P = 2000*(1.033)^x b) graph c) x = 8, P = 2000(1.033)^8

OpenStudy (anonymous):

oh my god... stupid error P (in millions) = 4(1.033)^x

OpenStudy (anonymous):

in 2008, x = 8: P = 4(1.033)^8 million